# On an Optimal Stopping Problem with a Discontinuous Reward

### v2

@@ 3.1. Surrender Right - Trivial Case
In this section, we derive a simple condition on the fee and surrender charge functions under which it is always optimal for a policyholder maximizing the risk-neutral value of her policy to hold the contract until maturity.
@@
Let $g$ and $\varphi$ be as defined as in (6). If the discounted surrender value process $Y=\{Y_{t}\}_{0\leq t\leq T}$, with $Y_{t}:=e^{-rt}F_{t}g(t,\,F_{t})$, is a submartingale, then the discounted reward process $Z=\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}:=e^{-rt}\varphi(t,F_{t})$, is also a submartingale.
In the proof below, and in the rest of the paper when necessary, we use the notation $\mathbb{E}_{s}[\cdot]=\mathbb{E}[\cdot|\mathcal{F}_{s}]$ for $s\leq t$.
@@
We first observe that for $0\leq t<T$, $Y_{t}=Z_{t}$ so that if $Y$ is a submartinale, then for any $0\leq s\leq t<T$, $\mathbb{E}_{s}[Z_{t}]\geq Z_{s}$. For $0\leq s\leq t=T$, since $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$, we have
@@
The next result is well-known in optimal stopping theory (see for example Björk (2009), Proposition 21.2) and is reproduced here for completeness. It states that if the discounted reward process is a submartingale, the maturity date of the contract is an optimal stopping time. There is a simple financial interpretation to the above statement. If the discounted reward process $Z$ is a submartingale, then it is non-decreasing on average over time. Hence, it is optimal to hold the contract as long as possible because its value is expected to increase. This is also the reasoning behind the optimal exercise strategy for an American call option on a non-dividend-paying stock.
@@
In the following, we denote $g_{x}(t,x)=\frac{\partial g}{\partial x}(t,x)$, $g_{xx}(t,x)=\frac{\partial^{2}g}{\partial x^{2}}(t,x)$ and $g_{t}(t,x)=\frac{\partial g}{\partial t}(t,x)$.
@@
Let $C$ and $g$ be the fee and the surrender charge functions as defined in (2) and (6), respectively. If
(9) $g_{t}(t,\,x)+(r-C(t,\,x)+\sigma^{2})xg_{x}(t,\,x)+\frac{\sigma^{2}x^{2}}{2}g_{xx}(t,\,x)-C(t,x)g(t,\,x)\geq 0,$
holds for all $(t,\,x)\in[0,T)\times\mathbb{R}_{+}$, then $T$ is an optimal stopping time for (7).
@@
for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, then $Y$ is a submartingale and so is the discounted reward process by Lemma 3.1. The optimal stopping time follows from Lemma 3.2. ∎
@@
In the next section, Corollary 4.3 shows that when the inequality in (9) is satisfied for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, the optimal stopping time $T$ is unique.
We now present applications of Proposition 3.1.
@@
Let the fee rate be constant, that is $C(t,x)=c>0$, and let ${g(t,x)=e^{-k(T-t)}}$ for some $k>0$. To eliminate the incentive to surrender, it suffices to choose $k$ and $c$ such that (9) holds; a quick calculation yields $k\geq c$. Therefore, it suffices to set the surrender rate $k$ at or above the fee rate $c$ to eliminate surrender risk. This result is well-known and discussed in Bernard and MacKay (2015), Proposition 3.1 and MacKay (2014), Proposition 4.4.2. In particular, MacKay (2014) shows that in the Black-Scholes setting, $g(t,x)=e^{-c(T-t)}$ is the maximal surrender charge function that can be used to eliminate the surrender incentive when the fee rate is constant.
@@
Let $g(t,x)=1-k\left(1-\frac{t}{T}\right)^{3}$ for some constant $k>0$, as in Bacinello and Zoccolan (2019), MacKay (2014) Section 4.3.2., and MacKay et al. (2017). As explained in MacKay (2014), this form of surrender charge function mimics the surrender charges in the market, which are usually high in the first years of the contract and drop drastically thereafter. Assume also that the fee function only depends on time, such that $C(t,\cdot)=c(t)$. In this example, we do not specify a particular form for the fee function; (9) is used to obtain the function $c(t)$ eliminating the surrender incentive. Simple calculations show that (9) is satisfied if
@@
Thus, any function $c(t)$ satisfying this inequality above will eliminate the surrender incentive (from a risk-neutral value maximization perspective). We may simply set $c(t)=\frac{3k/T\left(1-t/T\right)^{2}}{1-k(1-t/T)^{3}}$. This example highlights the interplay between the fee and the surrender charge structures. Under this setting, it would have been impossible to satisfy (9) using a constant fee rate, except with $c=0$, as the function on the right-hand side of (10) vanishes at maturity. Having a zero fee rate throughout the life of the VA contract is not a feasible solution from the insurer’s perspective since the guarantee at maturity must be financed. On the other hand, isolating $k$ in (10) gives
@@
We observe that the term on the right-hand side of the inequality goes to infinity as $t$ approaches $T$ if $c(t)$ is kept constant over time. This is obviously not a feasible solution either.
The examples presented above illustrate the interplay between fees, surrender charges, and surrender incentives. In order to eliminate early surrender incentives, surrender charges and fees must be structured conjointly; that is, a time-dependent fee structure should be paired with a time-dependent surrender charge to satisfy (9).
@@ 3.2. Surrender Right - Non-Trivial Case
In this section, we study the existence of an optimal stopping time for the variable annuity contract when (9) is not satisfied. To do so, the definitions below are required.
@@
    If, in addition, the trajectories of $Z$ are upper semi-continuous, then
@@
(12) $\tau^{x}_{t}=\inf\left\{t\leq s\leq T,\big{|}\,v(s,F_{s}^{t,\,x})=\varphi(s,F_{s}^{t,\,x})\right\}.$
@@
The discounted reward procevss $Z=\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}=e^{-rt}\varphi(\tau,F_{t})$, is positive and adapted with upper semi-continuous trajectories. Moreover, it is dominated by an integrable non-negative random variable, $\sup_{0\leq t\leq T}e^{-rt}\varphi(t,F_{t})$, so it is of class (D). The result then follows from Theorem 3.3.
The next theorem, inspired by Palczewski and Stettner (2010), provides an alternative representation for the optimal stopping problem in (7) and will be essential to prove the continuity of the value function.
Consider the set of time points ${\mathcal{H}_{N}:=\{t_{0},t_{1},\ldots,t_{N}\}}$ with $t_{n}=n\Delta_{N}$, where $\Delta_{N}:=T/N$, ${N\in\mathbb{N}^{+}}$. Like in option pricing, we use the qualifier “Bermudan” to describe a variable annuity contract that can only be surrendered at times in $\mathcal{H}_{N}$. The time-$t$ risk-neutral value of such a Bermudan contract is given by
(13) $b_{N}(t,x)=\sup_{\tau\in\mathcal{T}_{\Delta_{N}}\cap[t,T]}\mathbb{E}\left[e^{-r(\tau-t)}\varphi(\tau,F_{\tau}^{t,x})\right],$
where $\mathcal{T}_{\Delta_{N}}$ is the set of stopping times taking values in $\mathcal{H}_{N}$. On the other hand, an American variable annuity contract can be surrendered at any time in $[0,T]$ and has the value defined in (7). We expect the value of the Bermudan contract to converge to its American counterpart as $N\rightarrow\infty$. This idea is formalized below.
Lemma 3.3 (MacKay et al. (2023), Proposition 4.5).
Let $v$ and $b_{N}$ be defined by (7) and (13), respectively. Then, for any $(t,x)\in[0,T)\times\mathbb{R}_{+}$,
$v(t,x)=\lim_{N\rightarrow\infty}b_{N}(t,x).$
The proof of Lemma 3.3 can be found in MacKay et al. (2023).
Remark 3.3.
Lemma 3.3 can be extended to any non-negative adapted càdlàg reward process, without need to modify the arguments provided in the proof of Proposition 4.5 of MacKay et al. (2023).
We can now show that the value function (7) can be expressed in terms of a continuous reward function.
@@
(14) $v(t,x)=\sup_{\tau\in\mathcal{T}_{t,\,T}}\mathbb{E}\left[e^{-r(\tau-t)}\left[g(\tau,F_{\tau}^{t,x})F_{\tau}^{t,x}\vee h(\tau,\,F_{\tau}^{t,x})\right]\right],$
@@
$h(s,\,x)=\mathbb{E}\left[e^{-r(T-s)}\max(G,F_{T}^{s,x})\right],\quad s\in[0,T].$
@@
Let us consider a discrete, or Bermudan, version of the optimal stopping problem with $\mathcal{H}_{N}=\{t_{0},t_{1},\ldots,t_{N}\}$, $t_{n}=nT/N=n\Delta_{N}$. Define
$\widetilde{J}^{N}_{n}=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{\Delta_{N}}\cap[t_{n},T]}\mathbb{E}\left[e^{-r\tau}\left[g(\tau,F_{\tau})F_{\tau}\vee h(\tau,\,F_{\tau})\right]\Big{|}\mathcal{F}_{t_{n}}\right],\quad 0\leq n\leq N.$
Then $\widetilde{J}^{N}=\{\widetilde{J}_{n}^{N}\}_{n=0,1,2,\ldots,N}$ is the Snell envelope of the discrete discounted reward process $\{e^{-rt_{n}}(g(t_{n},F_{t_{n}})F_{t_{n}}\wedge h(t_{n},F_{t_{n}}))\}_{n=0}^{N-1}$ defined on $\mathcal{H}_{N}$ which admits the representation (see Lamberton (1998), Theorem 10.1.3)
(15) $\small\begin{array}[]{lll}\widetilde{J}^{N}_{N}&=e^{-rT}\left[g(T,F_{T})F_{T}\vee h(T,F_{T})\right],&\\ \widetilde{J}^{N}_{n}&=\max\left(e^{-rt_{n}}\left[g(t_{n},F_{t_{n}})F_{t_{n}}\vee h(t_{n},F_{t_{n}})\right],\mathbb{E}\left[\widetilde{J}^{N}_{n+1}|\mathcal{F}_{t_{n}}\right]\right),&\textrm{ for }0\leq n\leq N-1.\end{array}$
Similarly, using the dynamic programming principle, the Snell envelope of the Bermudan contract of the original reward function can be written as follows
(16) $\small\begin{array}[]{lll}J^{N}_{N}&=e^{-rT}\max(G,\,F_{T}),&\\ J^{N}_{n}&=\max\left(e^{-rt_{n}}g(t_{n},F_{t_{n}})F_{t_{n}},\mathbb{E}\left[J^{N}_{n+1}|\mathcal{F}_{t_{n}}\right]\right),&\textrm{for }0\leq n\leq N-1.\end{array}$
[…]
$\widetilde{J}^{N}_{n}$ $=\max\left(e^{-rt_{n}}\left[g(t_{n},F_{t_{n}})F_{t_{n}}\vee h(t_{n},F_{t_{n}})\right],\mathbb{E}\left[\widetilde{J}^{N}_{n+1}|\mathcal{F}_{t_{n}}\right]\right)$
$=\max\left(e^{-rt_{n}}g(t_{n},F_{t_{n}})F_{t_{n}}\vee\mathbb{E}[\mathbb{E}[e^{-r{T}}\max(G,F_{T})|\mathcal{F}_{t_{n+1}}]|\mathcal{F}_{t_{n}}],\mathbb{E}\left[\widetilde{J}^{N}_{n+1}|\mathcal{F}_{t_{n}}\right]\right)$
$=\max\left(e^{-rt_{n}}g(t_{n},F_{t_{n}})F_{t_{n}},\mathbb{E}[e^{-rt_{n+1}}h(t_{n+1},F_{t_{n+1}})|\mathcal{F}_{t_{n}}],\mathbb{E}\left[\widetilde{J}^{N}_{n+1}|\mathcal{F}_{t_{n}}\right]\right)$
$=\max\left(e^{-rt_{n}}g(t_{n},F_{t_{n}})F_{t_{n}},\mathbb{E}\left[\widetilde{J}^{N}_{n+1}|\mathcal{F}_{t_{n}}\right]\right)$
[…]
$=J^{N}_{n}.$
The third to last equality comes from $\widetilde{J}^{N}_{n}\geq e^{-rt_{n}}h(t_{n},F_{t_{n}})$ for any $n\in\{0,1,\dots,N\}$, by definition of $\widetilde{J}^{N}_{n}$. The final assertion follows from Lemma 3.3 and Remark 3.3, which states that the Bermudan VA contract converges to its American version as $N\rightarrow\infty$. ∎
@@
$v(t,x)$ $=\sup_{\tau\in\mathcal{T}_{t,\,T}}\mathbb{E}\left[e^{-r(\tau-t)}\left[g(\tau,F_{\tau}^{t,x})F_{\tau}^{t,x}\vee h(\tau,\,F_{\tau}^{t,x})\right]\right]$
$=\mathbb{E}\left[e^{-r(\widetilde{\tau}-t)}\left[g(\widetilde{\tau},F_{\widetilde{\tau}}^{t,x})F_{\widetilde{\tau}}^{t,x}\vee h(\widetilde{\tau},\,F_{\widetilde{\tau}}^{t,x})\right]\right]$
@@
This proves that $\bar{\tau}^{x}_{t}$ is optimal for (7). ∎
We are now interested in whether, and under which conditions, $\bar{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s., where $\tau_{t}^{x}$ is the smallest optimal stopping time for the original problem defined in Corollary 3.1. Generally, we can conclude that $\widetilde{\tau}_{t}^{x}\leq\tau_{t}^{x}\leq\bar{\tau}_{t}^{x}$. This idea is formalized in the next lemma.
@@
For $0\leq t\leq T$, the optimal stopping times defined in (12), (17) and (18) satisfy
[…]
@@
[…]

### v3

@@ 3.1. Surrender Right - Trivial Case
In this section, we derive a simple condition on the fee and surrender charge functions under which it is always optimal for the risk-neutral policyholder to hold the contract until maturity.
@@
Let $g$ and $\varphi$ be defined as in (6). If the discounted surrender value process $Y=\{Y_{t}\}_{0\leq t\leq T}$, with $Y_{t}:=e^{-rt}F_{t}g(t,\,F_{t})$, is a submartingale, then the discounted reward process $Z=\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}:=e^{-rt}\varphi(t,F_{t})$, is also a submartingale.
In the proof and throughout the paper, we use the notation $\mathbb{E}_{t}[\cdot]=\mathbb{E}[\cdot|\mathcal{F}_{t}]$.
@@
We first observe that for $0\leq t<T$, $Y_{t}=Z_{t}$ so that if $Y$ is a submartinale, then for any $0\leq s\leq t<T$, $\mathbb{E}_{s}[Z_{t}]\geq Z_{s}$.
For $0\leq s<t=T$, since $g(T,x)=1$ for all $x\in\mathbb{R}_{+}$, we have
@@
The next result is well known in optimal stopping theory (see, for example, Björk (2009), Proposition 21.2) and is reproduced here for completeness. It states that if the discounted reward process is a submartingale, maturity of the contract is an optimal stopping time. A simple financial interpretation is that if the discounted reward process $Z$ is a submartingale, it is non-decreasing on average over time. Holding the contract as long as possible is then optimal because its value is expected to increase. This is also the reasoning behind the optimal exercise strategy for an American call option on a non-dividend-paying stock.
@@
Going forward, let $g_{x}(t,x)=\frac{\partial g}{\partial x}(t,x)$, $g_{xx}(t,x)=\frac{\partial^{2}g}{\partial x^{2}}(t,x)$ and $g_{t}(t,x)=\frac{\partial g}{\partial t}(t,x)$.
@@
Let $C$ and $g$ be the fee and the surrender charge functions as defined by (2) and (6), respectively. If
(9) $g_{t}(t,\,x)+(r-C(t,\,x)+\sigma^{2})xg_{x}(t,\,x)+\frac{\sigma^{2}x^{2}}{2}g_{xx}(t,\,x)-C(t,x)g(t,\,x)\geq 0$
for all $(t,\,x)\in[0,T)\times\mathbb{R}_{+}$, $T$ is an optimal stopping time for (7).
@@
for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, $Y$ is a submartingale and so is the discounted reward process by Lemma 3.1. The optimal stopping time follows from Lemma 3.2. ∎
@@
In Section 4.4, Theorem 4.3(ii) shows that when the inequality in (9) is satisfied for all $(t,x)\in[0,T)\times\mathbb{R}_{+}$, the optimal stopping time $T$ is unique.
Next, we present some applications of Proposition 3.1.
@@
Let the fee rate be constant, that is $C(t,x)=c>0$, and let ${g(t,x)=e^{-k(T-t)}}$ for some $k>0$. To eliminate the incentive to surrender, it suffices to choose $k$ and $c$ such that (9) holds; a quick calculation yields $k\geq c$. In this setting, it suffices to set the surrender rate $k$ at or above the fee rate $c$ to eliminate optimal early surrenders. This result is well-known and discussed in Bernard and MacKay (2015), Proposition 3.1 and MacKay (2014), Proposition 4.4.2. In particular, MacKay (2014) shows that in the Black-Scholes setting, $g(t,x)=e^{-c(T-t)}$ is the maximal surrender charge function that can be used to remove the surrender incentive when the fee rate is constant.
@@
Let $g(t,x)=1-k\left(1-\frac{t}{T}\right)^{3}$ for some constant $k>0$, as in Bacinello and Zoccolan (2019), MacKay (2014) Section 4.3.2., and MacKay et al. (2017). As explained in MacKay (2014), this form of surrender charge function mimics the surrender charges in the market, which are usually high in the first years of the contract and drop drastically thereafter. Assume also that the fee function depends only on time such that $C(t,\cdot)=c(t)$. In this example, we do not specify a particular form for the fee function; (9) is used to obtain the form $c(t)$ eliminating the surrender incentive. Simple calculations show that (9) is satisfied if
@@
Thus, any function $c(t)$ that satisfies this inequality above will eliminate the surrender incentive (from a risk-neutral perspective). We may simply set $c(t)=\frac{3k/T\left(1-t/T\right)^{2}}{1-k(1-t/T)^{3}}$. This example highlights the interplay between the fee and the surrender charge structures. In this setting, it would have been impossible to satisfy (9) using a constant fee rate, except with $c=0$, as the function on the right-hand side of (10) vanishes at maturity. Having a zero fee rate throughout the life of the VA contract is not a feasible solution from the insurer’s perspective since the maturity guarantee must be financed. On the other hand, isolating $k$ in (10) gives
@@
The right-hand side of the inequality goes to infinity as $t$ approaches $T$ if $c(t)$ is kept constant, which gives another reason to allow the fee to vary over time.
The examples presented above illustrate the interplay between fees, surrender charges, and surrender incentives. In order to eliminate early surrender incentives, surrender charges and fees must be structured conjointly; that is, a time-dependent fee must be paired with a time-dependent surrender charge to satisfy (9).
@@ 3.2. Surrender Right - Non-Trivial Case
In this section, we study the existence of an optimal stopping time for the variable annuity contract in the general case, including when condition (9) is not satisfied. We first recall two definitions from probability and optimal stopping theory.
@@
    If the trajectories of $Z$ are upper semi-continuous, then
@@
(12) $\tau^{x}_{t}=\inf\left\{t\leq s\leq T\big{|}\,v(s,F_{s}^{t,\,x})=\varphi(s,F_{s}^{t,\,x})\right\}.$
@@
The discounted reward process $Z=\{Z_{t}\}_{0\leq t\leq T}$, with $Z_{t}=e^{-rt}\varphi(\tau,F_{t})$, is positive and adapted with upper semi-continuous trajectories. Moreover, it is dominated by an integrable non-negative random variable, $\sup_{0\leq t\leq T}e^{-rt}\varphi(t,F_{t})$, so it is of class (D). The result then follows from Theorem 3.3. ∎
Theorem 3.4, inspired by Palczewski and Stettner (2010) and presented below, provides an alternative representation for the optimal stopping problem in (7) and will be essential to prove the continuity of the value function.
@@
(13) $v(t,x)=\sup_{\tau\in\mathcal{T}_{t,\,T}}\mathbb{E}\left[e^{-r(\tau-t)}\left(g(\tau,F_{\tau}^{t,x})F_{\tau}^{t,x}\vee h(\tau,\,F_{\tau}^{t,x})\right)\right],$
@@
$h(t,\,x)=\mathbb{E}\left[e^{-r(T-t)}\max(G,F_{T}^{t,x})\right],\quad t\in[0,T].$
@@
We define the Snell envelope of the discounted continuous reward process $\{\widetilde{J}_{s}\}_{0\leq s\leq T}$ by
$\widetilde{J}_{s}=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,T}}\mathbb{E}[{e^{-r\tau}g(\tau,F_{\tau})F_{\tau}\vee h(\tau,F_{\tau})\mid\mathcal{F}_{s}}]$
for $0\leq s\leq T$, and recall that the Snell envelope of the discounted original reward process $\{J_{s}\}_{0\leq s\leq T}$ is defined by
$J_{s}=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,T}}\mathbb{E}\left[e^{-r\tau}\varphi(\tau,F_{\tau})\mid\mathcal{F}_{s}\right].$
We show that $\widetilde{J}_{s}=J_{s}$ a.s. for all $0\leq s\leq T$.
First observe that since $g(t,x)x\vee h(t,x)\geq\varphi(t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$, $\widetilde{J}_{s}\geq J_{s}$ for all $0\leq s\leq T$.
Now we show that $\widetilde{J}_{s}\leq J_{s}$ for all $0\leq s\leq T$. By definition of $v$, $g(t,x)x\vee h(t,x)\leq v(t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}_{+}$. It follows that
$\widetilde{J}_{s}\leq\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,T}}\mathbb{E}\left[e^{-r\tau}v(\tau,F_{\tau})\mid\mathcal{F}_{s}\right]=\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,T}}\mathbb{E}\left[J_{\tau}\mid\mathcal{F}_{s}\right].$
Since $\{J_{s}\}_{0\leq s\leq T}$ is a supermartingale, by the optimal sampling theorem, $\mathbb{E}[J_{\tau}\mid\mathcal{F}_{s}]\leq J_{s}$ for any $\tau\in\mathcal{T}_{s,T}$, and thus $\operatorname*{ess\,sup}_{\tau\in\mathcal{T}_{s,T}}\mathbb{E}\left[J_{\tau}\mid\mathcal{F}_{s}\right]\leq J_{s}$, which completes the proof.
@@
$v(t,x)$ $=\mathbb{E}\left[e^{-r(\widetilde{\tau}-t)}\left(g(\widetilde{\tau},F_{\widetilde{\tau}}^{t,x})F_{\widetilde{\tau}}^{t,x}\vee h(\widetilde{\tau},\,F_{\widetilde{\tau}}^{t,x})\right)\right]$
@@
It is possible to show that $\bar{\tau}_{t}^{x}=\tau_{t}^{x}$ a.s., where $\tau_{t}^{x}$ is the smallest optimal stopping time for the original problem defined in Corollary 3.1. A comparison of the three stopping time is given in lemma 3.3 below.
@@
For $0\leq t\leq T$, the optimal stopping times defined in (12), (14) and (15) satisfy $\widetilde{\tau}_{t}^{x}\leq\tau_{t}^{x}$ a.s. and $\tau_{t}^{x}=\bar{\tau}_{t}^{x}$ a.s.
@@
For $t=T$, the result trivially holds with equality. Henceforth, let $t<T$. To show $\widetilde{\tau}_{t}^{x}\leq\tau_{t}^{x}$, let $F_{t}=x$ and consider three cases.
(1) Suppose ${v(t,x)>g(t,x)x\vee h(t,x)\geq\varphi(t,x)}$. Then, for any $s\in[t,T]$, ${g(s,F_{s}^{t,x})F_{s}^{t,x}\vee h(s,\,F_{s}^{t,x})\geq\varphi(s,F_{s}^{t,x})}$ a.s., so, the process $v(s,F_{s})$ must first cross the continuous reward process $g(s,F_{s})F_{s}\vee h(s,F_{s})$ to attain the discontinuous reward process $\varphi(s,F_{s})$ (since the process $v(s,F_{s})$ starts above the two reward processes at $t$). It follows that $\widetilde{\tau}_{t}^{x}\leq\tau_{t}^{x}$. (2) If $v(t,x)=h(t,x)$ then necessarily $h(t,x)\geq xg(t,x)$, since by the continuous reward representation $v(t,x)\geq h(t,x)\vee xg(t,x)$, so that $\widetilde{\tau}_{t}^{x}=t$, and $\tau_{t}^{x}\geq t$, and the first inequality holds. The last case (3) is when $v(t,x)=xg(t,x)$, which automatically implies $\widetilde{\tau}_{t}^{x}=\tau_{t}^{x}=t$.
To show that $\tau_{t}^{x}=\bar{\tau}_{t}^{x}$, we fix $\omega\in\Omega$ and consider two cases: (1) $\tau_{t}^{x}(\omega)<T$ and (2) $\tau_{t}^{x}(\omega)=T$.
(1) By the definition of $\tau_{t}^{x}$ and the reward function $\varphi$, we have ${v(\tau_{t}^{x}(\omega),F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega))=g(\tau_{t}^{x}(\omega),F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega))F^{t,x}_{\tau_{t}^{x}(\omega)}(\omega)}$, and if $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)$, it follows that ${\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)=\bar{\tau}_{t}^{x}(\omega).}$ From the definition of $\tau_{t}^{x}(\omega)$, we know that ${v(s,F^{t,x}_{s}(\omega))>g(s,F^{t,x}_{s}(\omega))F^{t,x}_{s}(\omega)}$ for all ${t\leq s\leq\tau^{x}_{t}(\omega).}$ Thus, for $\widetilde{\tau}_{t}^{x}(\omega)$ to be equal to $\tau_{t}^{x}(\omega)$, we need ${v(s,F^{t,x}_{s}(\omega))>h(s,F^{t,x}_{s}(\omega))}$, for all ${t\leq s\leq\tau^{x}_{t}(\omega)}.$ Now, suppose there exists some $s$ in the interval $[t,\tau_{t}^{x}(\omega)]$ such that ${v(s,F^{t,x}_{s}(\omega))=h(s,F^{t,x}_{s}(\omega))}.$ In this case, the two stopping times $\tau_{t}^{x}(\omega)$ and $\bar{\tau}_{t}^{x}(\omega)$ would differ. However, the equality ${v(s,F^{t,x}_{s}(\omega))=h(s,F^{t,x}_{s}(\omega))}$ would imply that the probability that the contract is exercised before maturity is nil, which contradicts our initial assumption that $\tau_{t}^{x}(\omega)<T$. Therefore, it must hold that ${v(s,F^{t,x}_{s}(\omega))>h(s,F^{t,x}_{s}(\omega))}$ for all ${t\leq s\leq\tau^{x}_{t}(\omega)}$, which leads to $\widetilde{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)=\bar{\tau}_{t}^{x}(\omega).$
(2) $\tau_{t}^{x}(\omega)=T$ implies $v(s,F^{t,x}_{s}(\omega))>g(s,F^{t,x}_{s}(\omega))F^{t,x}_{s}(\omega)$ for all $t\leq s\leq T$. This means that if $\widetilde{\tau}_{t}^{x}(\omega)<T$, $v(\widetilde{\tau}_{t}^{x}(\omega),F_{\widetilde{\tau}_{t}^{x}(\omega)}(\omega))=h(\widetilde{\tau}_{t}^{x}(\omega),F_{\widetilde{\tau}_{t}^{x}(\omega)}(\omega))$ and by definition of $\bar{\tau}_{t}^{x}$, $\bar{\tau}_{t}^{x}(\omega)=\tau_{t}^{x}(\omega)$. ∎
The intuition behind the equality between $\tau^{x}_{t}$ and $\bar{\tau}^{x}_{t}$ is that $\bar{\tau}^{x}_{t}$ depends on why the contract is surrendered in the problem with the continuous reward. Indeed, the strategy underlying $\widetilde{\tau}^{x}_{t}$ is to stop if the contract value is equal to either the surrender benefit or the expected present value of the maturity benefit. If the reason to stop is equality with the surrender benefit, then this optimal stopping time coincides with the strategy underlying the problem with the discontinuous reward function and all three stopping times are equal. However, if the reason to stop is that the value of the contract is equal to the EPV of the maturity benefit, then $\bar{\tau}^{x}_{t}$ is set to maturity $T$. In that case, it is also optimal to wait until maturity in the original problem, since the contract being worth the EPV of the maturity benefit means that it is never be optimal to surrender before maturity.